# Life at low Reynolds number isn’t such a drag

### v1

@@ Abstract
We also learn as undergraduates that fluid flows at low Reynolds number are laminar. Despite living at low Reynolds number, dense collectives of actively moving microswimmers can overturn this paradigm. A prominent example is the phenomenon of “active turbulence”. Because many microswimmers (like bacteria, sperm, and even motor-driven biopolymer assemblies) are elongated in shape, they tend to align when close enough to each other—just like the molecules in a nematic liquid crystal. [^16] D. L. Koch & G. Subramanian, “Collective hydrodynamics of swimming microorganisms: living fluids”, Annual Review of Fluid Mechanics, 43, 637 (2011). Deviations from this alignment between neighbors are penalized, as described by a Frank elastic constant $K$. As a result, microswimmers in a dense collective often move in coherent swirls that, as they interact, give rise to large-scale chaotic flows reminiscent of turbulence! Ongoing research continues to explore the similarities and differences between this intriguing phenomenon and typical inertial turbulence. For example, just as the transition from laminar to inertial turbulence can be parameterized using the Reynolds number, recent work hints that a distinct dimensionless parameter — the “activity number” $A$ — may similarly described the transition to active turbulence. [^17] R. Alert, J. Casademunt, J. F. Joanny, “Active turbulence”, Annual Review of Condensed Matter Physics, 13, 143 (2022). One way to define this parameter is as $A=L_{s}/\sqrt{K/\zeta}$, which compares the system size $L_{s}$ to the length scale $\sqrt{K/\zeta}$above which active swimming overcomes the restoring force of nematic elasticity; here, $\zeta$ is an active stress coefficient, which can be thought of as being $\sim\Pi^{\mathrm{swim}}$, for example. In addition to unraveling these fundamental puzzles, an important direction for chemical engineering research is to develop ways to control active turbulence and harness it for e.g., improved solute mixing.
@@
In closing, I note that our discussion thus far has focused on the fascinating behaviors of microbes swimming in bulk fluids at low Reynolds number. But in many cases, microbes instead swim through elastic solids, such as the extracellular polymer matrix in a bacterial biofilm. Not only does confinement by the elastic matrix change how microbes swim, [^18] T. Bhattacharjee & S. S. Datta, “Bacterial hopping and trapping in porous media”, Nature Communications 10, 2075 (2019)., [^19] S. E. Spagnolie and P. T. Underhill, “Swimming in complex fluids”, Annual Review of Condensed Matter Physics, 14, 381 (2023). but the active stress generated by swimming can deform the matrix itself—unexpectedly leading to formation of self-sustained elastic standing waves. [^20] H. Xu, Y. Huang, R. Zhang, Y. Wu, “Autonomous waves and global motion modes in living active solids”, Nature Physics, 19, 46 (2023). One might imagine that the dimensionless parameter $\Pi^{\mathrm{swim}}/G$, where $G$ is the shear modulus of the elastic matrix, characterizes the onset of this behavior, although such a parameter has not, to my knowledge, yet been identified. Investigating the unusual properties of such “active solids” continues to be a fascinating direction for ongoing research.

### v2

@@ Abstract
We also learn as undergraduates that fluid flows at low Reynolds number are laminar. Despite living at low Reynolds number, dense collectives of actively moving microswimmers can overturn this paradigm. A prominent example is the phenomenon of “active turbulence”. Because many microswimmers (like bacteria, sperm, and even motor-driven biopolymer assemblies) are elongated in shape, they tend to align when close enough to each other—just like the molecules in a nematic liquid crystal. [^16] D. L. Koch & G. Subramanian, “Collective hydrodynamics of swimming microorganisms: living fluids”, Annual Review of Fluid Mechanics, 43, 637 (2011). Deviations from this alignment between neighbors are penalized, as described by a Frank elastic constant $K$. As a result, microswimmers in a dense collective often move in coherent swirls that, as they interact, give rise to large-scale chaotic flows reminiscent of turbulence! Ongoing research continues to explore the similarities and differences between this intriguing phenomenon and typical inertial turbulence. For example, just as the transition from laminar to inertial turbulence can be parameterized using the Reynolds number, recent work hints that a distinct dimensionless parameter — the “activity number” $A$ — may similarly described the transition to active turbulence. [^17] R. Alert, J. Casademunt, J. F. Joanny, “Active turbulence”, Annual Review of Condensed Matter Physics, 13, 143 (2022). One way to define this parameter is as $A=L_{s}/\sqrt{K/\zeta}$, which compares the system size $L_{s}$ to the length scale $\sqrt{K/\zeta}$above which active swimming overcomes the restoring force of nematic elasticity; here, $\zeta$ is an active stress coefficient, which can be thought of as being $\sim\Pi^{\mathrm{swim}}$, for example. [^18] A related parameter is the Ericksen number, $\mathrm{Er}=\eta VL/K$, which compares the force generated by swimming $\eta VL$ to the elastic restoring force $K$. In addition to unraveling these fundamental puzzles, an important direction for chemical engineering research is to develop ways to control active turbulence and harness it for e.g., improved solute mixing.
@@
In closing, I note that our discussion thus far has focused on the fascinating behaviors of microbes swimming in bulk fluids at low Reynolds number. But in many cases, microbes instead swim through elastic solids, such as the extracellular polymer matrix in a bacterial biofilm. In this case, confinement by the elastic matrix changes how microbes swim. [^19] T. Bhattacharjee & S. S. Datta, “Bacterial hopping and trapping in porous media”, Nature Communications 10, 2075 (2019). This change can be described by two new dimensionless parameters: the Deborah number $\mathrm{De}=t_{\mathrm{mat}}/t_{\mathrm{swim}}$ comparing the relaxation time scale of the matrix $t_{\mathrm{mat}}$ to a characteristic time scale of swimming $t_{\mathrm{swim}}$, and the Benes number $\mathrm{Bn}=L_{\mathrm{mat}}/L$ comparing the confinement length scale imposed by the matrix $L_{\mathrm{mat}}$ to the size of the microbe $L$ [^20] S. E. Spagnolie and P. T. Underhill, “Swimming in complex fluids”, Annual Review of Condensed Matter Physics, 14, 381 (2023). But also, the active stress generated by swimming can deform the matrix itself—unexpectedly leading to formation of self-sustained elastic standing waves. [^21] H. Xu, Y. Huang, R. Zhang, Y. Wu, “Autonomous waves and global motion modes in living active solids”, Nature Physics, 19, 46 (2023). One might imagine that the dimensionless parameter $\Pi^{\mathrm{swim}}/G$, where $G$ is the shear modulus of the elastic matrix, characterizes the onset of this behavior, although such a parameter has not, to my knowledge, yet been identified. Investigating the unusual properties of such “active solids” continues to be a fascinating direction for ongoing research.